### v2

# A Recursive Theory of Variational State Estimation: The Dynamic Programming Approach

Abstract
In this article, variational state estimation is examined from the dynamic programming perspective. This leads to two different value functional recursions depending on whether backward or forward dynamic programming is employed. The result is a theory of variational state estimation that corresponds to the classical theory of Bayesian state estimation. More specifically, in the backward method, the value functional corresponds to a likelihood that is upper bounded by the state likelihood  [^1] This is often referred to as the “backward filter”. from the Bayesian backward recursion. In the forward method, the value functional corresponds to an unnormalized density that is upper bounded by the unnormalized filtering density. Both methods can be combined to arrive at a variational two-filter formula. Additionally, it is noted that optimal variational filtering is generally of quadratic time-complexity in the sequence length. This motivates the notion of sub-optimal variational filtering, which also lower bounds the evidence but is of linear time-complexity. Another problem is the fact that the value functional recursions are generally intractable. This is briefly discussed and a simple approximation is suggested that retrieves the filter proposed by Courts2021. The methodology is examined in (i) a jump Gauss–Markov system under a certain factored Markov process approximation, and (ii) in a Gauss–Markov model with log-polynomial likelihoods under a Gauss–Markov constraint on the variational approximation. It is demonstrated that the value functional recursions are tractable in both cases. The resulting estimators are examined in simulation studies and are found to be of adequate quality in comparison to sensible baselines.
1 Introduction
Consider a partially observed Markov process given by
$x_{0}$ $\sim\pi_{0}(x_{0})$ (1a)
$x_{t}\mid x_{t-1}$ $\sim\pi_{t\mid t-1}(x_{t}\mid x_{t-1})$ (1b)
$y_{t}\mid x_{t}$ $\sim h_{t\mid t}(y_{t}\mid x_{t})$ (1c)
on the discrete interval $0\leq t\leq T$ for some fixed $T$. Here $y_{0\mathrel{\mathop{\ordinarycolon}}T}$ is the observed sequence, $x_{0\mathrel{\mathop{\ordinarycolon}}T}$ is the (hidden) state sequence, $\pi_{0}$ is the initial state density, $\pi_{t\mid t-1}$ are the transition kernels, and $h_{t\mid t}$ is are observation kernels. The problem of state estimation is to infer the corresponding state sequence $x_{0\mathrel{\mathop{\ordinarycolon}}T}$. This kind of estimation problem commonly arise in science and engineering, such as in tracking and navigation BarShalom2004, TittertonWeston2004, MarkleyCrassidis2014, time series analysis (Durbin2012, Hyndman2008), finance (Lindstrom2018), robotics (Barfoot2017), geoscience (Carrassi2018), data assimilation (Evensen2009), and neuroscience (Galka2004), to give a few examples.
From the Bayesian perspective, this entails computing the a posteriori distribution over state paths $x_{0\mathrel{\mathop{\ordinarycolon}}T}$ given the observed path $y_{0\mathrel{\mathop{\ordinarycolon}}T}$. Formally, the problem is well understood and can be solved by either a backward or forward recursion, the reader is referred to, for instance, Ho1964, Kitagawa1987, Cappe2005, Sarkka2023, a comprehensive account of the linear (Gauss–Markov) theory is given by Kailath2000.
However, in general, these recursions involve a series of intractable integrals. Sequential Monte Carlo methods provide a principled approach to approximating functionals with respect to the posterior over state paths (Gordon1993, Pitt1999, Arulampalam2002, Crisan2002, Djuric2003, Cappe2007, Kantas2009, Briers2010, Guarniero2017, Naesseth2019).
A popular and more classical approach to approximate to state estimation is the assumed density method (Maybeck1982), which “approximates as you go” within the Bayesian forward recursion (Bell1993, Julier2000, Lefebvre2002, Wu2006, Tronarp2019b, Corenflos2023, Sarkka2023). These approximations can then be refined via iterative schemes Bell1994, Garcia2015, Garcia2016, Tronarp2018, Garcia2019a. However, these approximations can be brittle and are not derived from some principle of optimality, which make them difficult to analyze.
On the other hand, variational inference allows for control of the class of approximating densities and is optimal in relative entropy sense (Beal2003, Blei2017). Thus the need to assume a density is obviated. While variational inference has long been used in state estimation, it has mostly been employed as a local approximation within the assumed density framework (Smidl2008, Darling2017, Gultekin2017, Zhao2020, Courts2021). Some approaches to global variational state estimation has been suggested by Agamennoni2012, Courts2021, but these lack the recursive structure akin to Bayesian state estimation theory. Results in this direction have instead been obtained by Weber2015, whom develop a backward recursion for the variational lower bound, see also Kim2020. Further, a forward recursion for the variational lower bound was developed by Campbell2021, for which error estimates have been obtained by (Chagneux2024).
While the approaches of Weber2015 and Campbell2021 are indeed recursive, their relation to classical Bayesian state estimation theory remains opaque. A central aim of this article is to develop a theory of recursive variational state estimation, which is completely analogous to Bayesian state estimation theory. Therefore, before stating the present contribution in full, a brief review Bayesian state estimation theory is given in section˜1.1, which also serves to fix some notational conventions used throughout. The present contribution is then outlined in section˜1.2.
1.1 Bayesian state estimation theory
There are essentially two ways to construct the a posteriori distribution over state paths, the forward recursion and the backward recursion (Cappe2005), see also Ho1964, Askar1981, Kitagawa1987. The result in a forward and reverse-time Markov factorization of the posterior, respectively. In order to state the result, introduce the following notation. The conditional distribution of $x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}$ given $x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}$ and $y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}$ be denoted by
$\pi_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}\mid t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}^{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}(x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}\mid x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}},y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}),$ (2)
where in particular $\pi_{t}^{1\mathrel{\mathop{\ordinarycolon}}t}$ are the filtering distributions and $\pi_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}$ are the smoothing marginals. Similarly, let the conditional distribution of $y_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}$ given $x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}$ and $y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}$ be denoted by
$h_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}\mid t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}^{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}(y_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}\mid x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}},y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}).$ (3)
For example, $h_{1\mathrel{\mathop{\ordinarycolon}}t}$ is the marginal likelihood associated with $y_{0\mathrel{\mathop{\ordinarycolon}}t}$ and $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$ is the likelihood function for $x_{t}$ given the future observations $y_{t+1\mathrel{\mathop{\ordinarycolon}}T}$. Also, formally define $h_{T+1\mathrel{\mathop{\ordinarycolon}}T\mid T}(x_{T})=1$. Furthermore, define the unnormalized conditional density of $x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}$ given $x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}$ and $y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}$ by
$\bar{\pi}_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}\mid t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}^{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}(x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}\mid x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}},y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}})=h_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}\mid t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}(y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}\mid x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}})\pi_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}\mid t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}^{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}(x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}\mid x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}).$ (4)
For example, $\bar{\pi}_{t\mid t-1}^{t}=h_{t\mid t}\pi_{t\mid t-1}$, the unnormalized filtering density satisfy $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}=h_{0\mathrel{\mathop{\ordinarycolon}}t}\pi_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$, and the unnormalized smoothing density satisfy $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}=h_{0\mathrel{\mathop{\ordinarycolon}}T}\pi_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}$. The backward method establishes a recursion over $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$ from which a forward Markov representation of $\pi_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}$ is obtained, this is theorem˜1.
Theorem 1 (Bayesian backward method).
The likelihood $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$ satisfies the following recursion
$h_{t\mathrel{\mathop{\ordinarycolon}}T\mid t-1}(x_{t-1})=\int h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}(x_{t})\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})\operatorname{d\!}x_{t}.$ (5)
Additionally, the marginal likelihood is given by
$h_{0\mathrel{\mathop{\ordinarycolon}}T}=\int h_{1\mathrel{\mathop{\ordinarycolon}}T\mid 0}(x_{0})\bar{\pi}_{0}^{0}(x_{0})\operatorname{d\!}x_{0},$ (6)
and the state path has the following a posteriori forward Markov representation
$\pi_{0}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0})$ $=\frac{h_{1\mathrel{\mathop{\ordinarycolon}}T\mid 0}(x_{0})\bar{\pi}_{0}^{0}(x_{0})}{h_{0\mathrel{\mathop{\ordinarycolon}}T}}$ (7a)
$\pi_{t\mid t-1}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{t}\mid x_{t-1})$ $=\frac{h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}(x_{t})\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})}{h_{t\mathrel{\mathop{\ordinarycolon}}T\mid t-1}(x_{t-1})}.$ (7b)
There is a also a corresponding forward method, the so-called filtering recursion. It is usually stated in terms of the the normalized filtering density, $\pi_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$. However, in the following, the result is instead stated in terms of a recursion over the unnormalized filtering density, $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$. The reason is that the unnormalized filtering density is easier to relate to the variational state estimation theory, which is to be developed. The filtering recursion and its relation to the posterior, $\pi_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}$, is stated in theorem˜2.
Theorem 2 (Bayesian forward method).
The unnormalized filtering density satisfies the following recursion
$\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})=\int\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})\bar{\pi}_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}(x_{t-1})\operatorname{d\!}x_{t-1}.$ (8)
Additionally, the marginal likelihood is given by
$h_{0\mathrel{\mathop{\ordinarycolon}}T}=\int\bar{\pi}_{T}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{T})\operatorname{d\!}x_{T},$ (9)
and the state path has the following a posteriori reverse-time Markov representation
$\pi_{T}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{T})$ $=\frac{\bar{\pi}_{T}^{T}(x_{T})}{h_{0\mathrel{\mathop{\ordinarycolon}}T}}$ (10a)
$\pi_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{t-1}\mid x_{t})$ $=\frac{\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})\bar{\pi}_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}(x_{t-1})}{\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})}.$ (10b)
Lastly, the likelihoods and filtering densities may be combined to obtain the smoothing marginals $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}$. This is the so-called “two-filter” formula (Bresler1986), which is given by corollary˜1.
Corollary 1 (The two-filter formula).
The unnormalized smoothing density, $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}$, satisfy the following relation
$\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{t})=h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}(x_{t})\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t}).$ (11)
1.2 Contribution and outline
In variational inference an appropriate class of densities, $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$, is selected and the posterior is approximated by the best one in the sense of the following lower bound on the marginal likelihood (Beal2003, Blei2017)
$\log h_{0\mathrel{\mathop{\ordinarycolon}}T}\geq\underset{q_{0\mathrel{\mathop{\ordinarycolon}}T}\in\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}}{\max}\,\int\log\Big\{\frac{\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})}{q_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})}\Big\}q_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})\operatorname{d\!}x_{0\mathrel{\mathop{\ordinarycolon}}T}$
Recursive approaches have previously been formulated by Weber2015 and Campbell2021 from the reinforcement learning perspective. Here an optimal control approach is taken instead where the time-marginals of $q_{0\mathrel{\mathop{\ordinarycolon}}T}$ are states and the associated transition kernels are the controls. The problem can then formally be solved by both backward and forward dynamic programming (Bellman1966, Larson1966), which involve recursions over associated value functionals.  [^2] It ought to be noted that the present notion of value functionals differs from the notion of value function used by Weber2015, Campbell2021. The relationship between the present contribution and that of Weber2015, Campbell2021 is further discussed in section 5. These methods shall be referred to as the variational backward and forward methods, respectively. Based on this a theory of recursive variational state estimation is developed in section˜2, which goes beyond the contributions of Weber2015, Campbell2021.
More specifically, a variational analogue of the Bayesian state estimation is obtained by establishing:
  • An analogue of theorem˜1 is obtained by the variational backward method. The associated value functional is completely determined by a likelihood $\beta_{t}$, which is upper bounded by $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$. This is established in theorem˜3 and corollary˜2.
  • An analogue of theorem˜2 is obtained by the variational forward method. The associated value functional is completely determined by an unnormalized density $\alpha_{t}$, which is upper bounded by $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$. This is established in theorem˜4 and corollary˜3.
  • An analogue of corollary˜1 is obtained by combining theorem˜3 and theorem˜4. More specifically, the time-marginals of the variational posterior are relative entropy optimal approximations of the normalization of $\alpha_{t}\beta_{t}$. This is corollary˜5.
The recursions of theorem˜3 and theorem˜2 are parametrized by the time-marginals of the variational posterior, which are of course unavailable. This problem is circumvented by the fixed-point equations of propositions˜1 and 2. Furthermore, the notion of variational filtering is discussed in section˜3. In particular, the preceeding analysis suggests that the optimal variational filter is in general of time-complexity $O(T^{2})$. This leads to the notion of sub-optimal variational filtering, which is of time-compelxity $O(T)$. Another problem is that the value functional recursions are intractable in general. This is briefly discussed in section˜4, where the fact that $\alpha_{t}$ is an unnormalized density is used together with variational inference to obtain an approximation. This procedure is shown to recover the filter proposed by Courts2021.
Additionally, two case studies are given where the value functional recursions are indeed tractable. In section˜6 jump-Gauss Markov systems are studied when the a posterior distribution of the state and jump processes is approximated as independendent Markov processes and in section˜7 a Gauss–Markov model with log-polynomial likelihoods is examined under a Gauss–Markov constraint.
2 Variational state estimation theory
Let $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ be a family of distributions over stochastic processes indexed by $0,\ldots,T$, and assume that $\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}$ absolutely continuous with respect to $q_{0\mathrel{\mathop{\ordinarycolon}}T}$ for each $q_{0\mathrel{\mathop{\ordinarycolon}}T}$ in $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$. Then for any $q_{0\mathrel{\mathop{\ordinarycolon}}T}\in\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$, the marginal likelihood, $h_{0\mathrel{\mathop{\ordinarycolon}}T}$, is given by
$h_{0\mathrel{\mathop{\ordinarycolon}}T}=\int\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})\operatorname{d\!}x_{0\mathrel{\mathop{\ordinarycolon}}T}=\int\frac{\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})}{q_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})}q_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})\operatorname{d\!}x_{0\mathrel{\mathop{\ordinarycolon}}T}.$ (12)
Furthermore, by Jensen’s inequality the logarithm of $h_{0\mathrel{\mathop{\ordinarycolon}}T}$ is bounded from below by
$\log h_{0\mathrel{\mathop{\ordinarycolon}}T}\geq\int\log\Big\{\frac{\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})}{q_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})}\Big\}q_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})\operatorname{d\!}x_{0\mathrel{\mathop{\ordinarycolon}}T}.$ (13)
This lower bound is referred to as evidence lower bound, variational lower bound, or free energy (Beal2003, Blei2017), which prompts the following definition.
Definition 1.
Given an unnormalized posterior, $\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}$, the variational lower bound functional, $\mathcal{V}$, is defined as
$\mathcal{V}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T},q_{0\mathrel{\mathop{\ordinarycolon}}T})=\int\log\Big\{\frac{\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})}{q_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})}\Big\}q(x_{0\mathrel{\mathop{\ordinarycolon}}T})\operatorname{d\!}x_{0\mathrel{\mathop{\ordinarycolon}}T}.$ (14)
The idea behind variational inference is then to find the maximizer $q_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}$ in $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ of $\mathcal{V}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T},q_{0\mathrel{\mathop{\ordinarycolon}}T})$ Beal2003. This gives the notion of variational posterior, whose formal definition is stated as follows.
Definition 2 (Variational posterior).
Given a variational lower bound, $\mathcal{V}_{0\mathrel{\mathop{\ordinarycolon}}T}$, and a family of distributions, $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$, the variational posterior is defined as
$q_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}=\underset{q_{0\mathrel{\mathop{\ordinarycolon}}T}\in\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}}{\arg\,\max}\,\mathcal{V}_{0\mathrel{\mathop{\ordinarycolon}}T}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T},q_{0\mathrel{\mathop{\ordinarycolon}}T}).$ (15)
It is well-known that the variational posterior is the optimal approximation to $\pi_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}$ in relative entropy sense (Beal2003, Blei2017).
2.1 Markov Factorization of the variational family
The problem is then to define a variational family for which the variational posterior is tractable, computationally efficient, and accurate. From section˜1.1, it is already known that the posterior, $\pi_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}$, has both a forward and reverse-time Markov representation. Consequently, it is sufficient to only consider variational families for which each element has a Markov factorization. Assume that the elements of the variational family, $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$, admit a forward Markov factorization:
$q_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})=q_{0}(x_{0})\prod_{t=1}^{T}q_{t\mid t-1}(x_{t}\mid x_{t-1}).$ (16)
From $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ classes of time-marginals, $\mathcal{Q}_{t}$, and conditional densities, $\mathcal{Q}_{t+1\mathrel{\mathop{\ordinarycolon}}s\mid t}$ for $s>t$, are defined in the obvious way. For instance $q_{t}\in\mathcal{Q}_{t}$ implies that there is a $q_{0\mathrel{\mathop{\ordinarycolon}}T}\in\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ such that
$q_{t}(x_{t})=\int q_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})\operatorname{d\!}x_{0\mathrel{\mathop{\ordinarycolon}}t-1,t+1T}.$
In particular this means if $q_{0\mathrel{\mathop{\ordinarycolon}}T}\in\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ then the factors of (16) lies in $\mathcal{Q}_{0}$ and $\mathcal{Q}_{t\mid t-1}$, respectively. Furthermore, (16) also has a reverse-time Markov representation. Namely, by Bayes’ rule
$q_{t-1}(x_{t-1})q_{t\mid t-1}(x_{t}\mid x_{t-1})=q_{t}(x_{t})q_{t-1\mid t}(x_{t-1}\mid x_{t}).$ (17)
Consequently, any element in $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ also admit a reverse-time Markov factorization
$q_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})=q_{T}(x_{T})\prod_{t=0}^{T-1}q_{t\mid t+1}(x_{t}\mid x_{t+1}).$ (18)
Define the classes of conditional densities, $\mathcal{Q}_{s\mathrel{\mathop{\ordinarycolon}}t\mid t+1}$ for $s<t$, defined in the analogous way to $\mathcal{Q}_{t+1\mathrel{\mathop{\ordinarycolon}}s\mid t}$, then the conditional densities in (18) lie in $\mathcal{Q}_{s\mathrel{\mathop{\ordinarycolon}}t\mid t+1}$ and $q_{T}$ lie in $\mathcal{Q}_{T}$. The forward and reverse-time Markov factorizations can also be combined like so
$q_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})=q_{0\mathrel{\mathop{\ordinarycolon}}t-1\mid t}(x_{0\mathrel{\mathop{\ordinarycolon}}t-1}\mid x_{t})q_{t}(x_{t})q_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}(x_{t+1\mathrel{\mathop{\ordinarycolon}}T}\mid x_{t}).$ (19)
With these Markov factorizations, efficient algorithms for obtaining the variational posterior can be obtained by means of dynamic programming (Bellman1966). The key to this programme is the following decomposition of the variational lower bound
$\mathcal{V}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T},q_{0\mathrel{\mathop{\ordinarycolon}}T})=\mathcal{V}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}t}^{0\mathrel{\mathop{\ordinarycolon}}t},q_{0\mathrel{\mathop{\ordinarycolon}}t})+\int\mathcal{V}(\bar{\pi}_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}^{t+1\mathrel{\mathop{\ordinarycolon}}T},q_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t})q_{t}(x_{t})\operatorname{d\!}x_{t\mathrel{\mathop{\ordinarycolon}}T}$ (20)
where the fact that $\mathcal{V}(\bar{\pi}_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}^{t+1\mathrel{\mathop{\ordinarycolon}}T},q_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t})$ is a function in $x_{t}$ is omitted from the notation. In section˜2.2, the forward Markov factorization (16) is employed together with backward dynamic programming, giving a variational anlogue of theorem˜1. In section˜2.3 the reverse-time Markov factorization (18) is employed together with forward dynamic programming, which gives a variational analogue of theorem˜2. Lastly, in section˜2.4 the mixed forward/reverse-time factorization (19) is employed together with both forward and backward dynamic programming to obtain a variational analogue of corollary˜1.
2.2 Forward Markov factorization of the variational family
In this section the factorization (16) is used in combination with backward dynamic programming, where the variational inference problem is formally solved by a backward recursion for a particular value function. This value functional is obtained by maximizing the second term of (20) with respect to $q_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$, which gives the notion of variational backward value functional.
Definition 3 (Variational backward value functional).
Define the variational backward value functional at time $T$ by $V^{\mathsf{b}}_{T}(q_{T})=0$ and for times $0\leq t<T$ by
$V^{\mathsf{b}}_{t}(q_{t})=\underset{q_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}\in\mathcal{Q}_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid T}}{\max}\,\int\mathcal{V}(\bar{\pi}_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}^{t+1\mathrel{\mathop{\ordinarycolon}}T},q_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t})q_{t}(x_{t})\operatorname{d\!}x_{t\mathrel{\mathop{\ordinarycolon}}T}.$
From definition˜3, the evidence lower bound satisfies the following decomposition
$\underset{q_{0\mathrel{\mathop{\ordinarycolon}}T}\in\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}}{\max\,}\mathcal{V}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T},q_{0\mathrel{\mathop{\ordinarycolon}}T})=\underset{q_{0\mathrel{\mathop{\ordinarycolon}}t}\in\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}t}}{\max\,}\Big[\mathcal{V}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}t\mid t}^{0\mathrel{\mathop{\ordinarycolon}}t},q_{0\mathrel{\mathop{\ordinarycolon}}t})+V^{\mathsf{b}}_{t}(q_{t})\Big].$
In particular, the initial distribution of the variational posterior may be obtained by
$q_{0}^{0\mathrel{\mathop{\ordinarycolon}}T}=\arg\,\max_{q_{0}\in\mathcal{Q}_{0}}\,\Big[\int\log\Big\{\frac{\bar{\pi}_{0}^{0}(x_{0})}{q_{0}(x_{0})}\Big\}q_{0}(x_{0})\operatorname{d\!}x_{0}+V^{\mathsf{b}}_{0}(q_{0})\Big].$ (21)
The variational backward value functional satisfies a particular recursion. More specifically, it is easy to see that
$\mathcal{V}(\bar{\pi}_{t\mathrel{\mathop{\ordinarycolon}}T\mid t-1}^{t\mathrel{\mathop{\ordinarycolon}}T},q_{t\mathrel{\mathop{\ordinarycolon}}T\mid t-1})=\mathcal{V}(\bar{\pi}_{t\mid t-1}^{t},q_{t\mid t-1})+\int\mathcal{V}(\bar{\pi}_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}^{t+1\mathrel{\mathop{\ordinarycolon}}T},q_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t})q_{t\mid t-1}(x_{t}\mid x_{t-1})\operatorname{d\!}x_{t}.$
Inserting this into definition˜3 then gives the variational backward value functional at time t1t-1 according to
$V^{\mathsf{b}}_{t-1}(q_{t-1})=\underset{q_{t\mid t-1}\in\mathcal{Q}_{t\mid t-1}}{\max\,}\Bigg[\int\mathcal{V}(\bar{\pi}_{t\mid t-1}^{t},q_{t\mid t-1})q_{t-1}(x_{t-1})\operatorname{d\!}x_{t-1}+V^{\mathsf{b}}_{t}(q_{t})\Bigg].$
The boundary constraint $V^{\mathsf{b}}_{T}(q_{T})=0$ of definition˜3 ensure that this recursion is valid for $0<t\leq T$. It is important to note that $q_{t}$ is obtained by marginalizing $q_{t\mid t-1}$ with respect to $q_{t-1}$. Respecting this constraint gives the following result.
Lemma 1.
The variational backward value functional satisfies the following recursion
$V^{\mathsf{b}}_{t-1}(q_{t-1})=\max_{q_{t\mid t-1}\in\mathcal{Q}_{t\mid t-1}}\,\Big[\int\mathcal{V}(\bar{\pi}_{t\mid t-1}^{t},q_{t\mid t-1})q_{t-1}(x_{t-1})\operatorname{d\!}x_{t-1}+V^{\mathsf{b}}_{t}(q_{t})\Big],$ (22)
where the maximization is subject to the following constraint
$q_{t}(x_{t})=\int q_{t\mid t-1}(x_{t}\mid x_{t-1})q_{t-1}(x_{t-1})\operatorname{d\!}x_{t-1}.$ (23)
The recursion of lemma˜1 is rather opaque but it turns out that the variational backward value functional is in fact a linear functional. Namely, the next step is to show that $V^{\mathsf{b}}_{t}$ admit the following representation
$V^{\mathsf{b}}_{t}(q_{t})=\int\log\{\beta_{t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}.$ (24)
As the backward representer $\beta_{t}$ completely determines the backward value functional $V^{\mathsf{b}}_{t}$, they are both referred to as the backward value functional at times. According to definition˜3, the terminal condition of $\beta$ must be $\beta_{T}(x_{T})=1$. Furthermore, if $q_{t\mid t-1}\in\mathcal{Q}_{t\mid t-1}$ and $q_{t}$ satisfies the constraint (23), then $V^{\mathsf{b}}_{t}$ is given by
$V^{\mathsf{b}}_{t}(q_{t})=\int\log\{\beta_{t}(x_{t})\}q_{t\mid t-1}(x_{t}\mid x_{t-1})q_{t-1}(x_{t-1})\operatorname{d\!}x_{t-1\mathrel{\mathop{\ordinarycolon}}t}.$
Inserting this into the recursion of lemma˜1, expanding $\mathcal{V}(\bar{\pi}_{t\mid t-1}^{t},q_{t\mid t-1})$ using definition˜1, and invoking the principle of induction gives the following result.
Theorem 3.
The variational backward value functional satisfies the following recursion
$\begin{split}&\int\log\{\beta_{t-1}(x_{t-1})\}q_{t-1}(x_{t-1})\operatorname{d\!}x_{t-1}\\ &=\underset{q_{t\mid t-1}\in\mathcal{Q}_{t\mid t-1}}{\max}\,\int\Big(\int\log\Big\{\frac{\beta_{t}(x_{t})\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})}{q_{t\mid t-1}(x_{t}\mid x_{t-1})}\Big\}q_{t\mid t-1}(x_{t}\mid x_{t-1})\operatorname{d\!}x_{t}\Big)q_{t-1}(x_{t-1})\operatorname{d\!}x_{t-1}.\end{split}$
Additionally, the initial density of the variational posterior satisfies
$q_{0}^{0\mathrel{\mathop{\ordinarycolon}}T}=\underset{q_{0}\in\mathcal{Q}_{0}}{\arg\,\max}\,\int\log\Big\{\frac{\beta_{0}(x_{0})\bar{\pi}_{0}^{0}(x_{0})}{q_{0}(x_{0})}\Big\}q_{0}(x_{0})\operatorname{d\!}x_{0},$
and the variational lower bound is the corresponding maximia.
Theorem˜3 allows for the recursive computation of a forward Markov factorization of the variational posterior in principle. However, there is a practical problem that the recursion for the representer $\beta_{t}$ depends on the time marginals $q_{0},\ldots,q_{T-1}$, which in fact need to be the time-marginals of the variational posterior. The solution is to alternative between computing value functionals and transition kernels and time-marginals, which lead to the fixed-point equations given by proposition˜1. In practice, the fixed-point iteration needs to be initialized. It appears difficult to construct an intelligent initializer based on the backward method but the forward method to be discussed in section˜2.3 can be initialized based on sub-optimal variational filtering to be discussed in section˜3.
Proposition 1 (Backward variational fixed-point equations).
The backward value functional and the variational posterior satisfy the following fixed-point equations for $1\leq t\leq T$
$q_{t\mid t-1}^{0\mathrel{\mathop{\ordinarycolon}}T}$ $=\underset{q_{t\mid t-1}\in\mathcal{Q}_{t\mid t-1}}{\arg\,\max}\,\int\Big(\int\log\Big\{\frac{\beta_{t}\bar{\pi}_{t\mid t-1}^{t}}{q_{t\mid t-1}}\Big\}q_{t\mid t-1}\operatorname{d\!}x_{t}\Big)q_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}T}\operatorname{d\!}x_{t-1}$ (25a)
$\log\beta_{t-1}$ $=\int\log\Big\{\frac{\beta_{t}\bar{\pi}_{t\mid t-1}^{t}}{q_{t\mid t-1}^{0\mathrel{\mathop{\ordinarycolon}}T}}\Big\}q_{t\mid t-1}^{0\mathrel{\mathop{\ordinarycolon}}T}\operatorname{d\!}x_{t}$ (25b)
$q_{0}^{0\mathrel{\mathop{\ordinarycolon}}T}$ $=\underset{q_{0}\in\mathcal{Q}_{0}}{\arg\,\max}\,\int\log\Big\{\frac{\beta_{0}\bar{\pi}_{0}^{0}}{q_{0}}\Big\}q_{0}\operatorname{d\!}x_{0}$ (25c)
$q_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}$ $=\int q_{t\mid t-1}^{0\mathrel{\mathop{\ordinarycolon}}T}q_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}T}\operatorname{d\!}x_{t-1}.$ (25d)
2.2.1 The backward representer as a likelihood function
Theorem˜3 suggest that the backward representers, $\beta_{t}$, serve a similar purpose in the variational state estimation as the likelihood $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$ serve in Bayesian state estimation (c.f. theorem˜1). In fact, the former is a lower bound to the latter since by Jensen’s inequality
$\begin{split}\log\beta_{t-1}(x_{t-1})&=\int\log\Big\{\frac{\beta_{t}(x_{t})\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})}{q_{t\mid t-1}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{t}\mid x_{t-1})}\Big\}q_{t\mid t-1}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{t}\mid x_{t-1})\operatorname{d\!}x_{t}\\ &\leq\log\int\beta_{t}(x_{t})\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})\operatorname{d\!}x_{t}.\end{split}$
Therefore, $\beta$, satisfies the following difference inequality
$\beta_{t-1}(x_{t-1})\leq\int\beta_{t}(x_{t})\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})\operatorname{d\!}x_{t},$ (26)
but $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$ satisfies this relation with strict equality. Thus if $\beta_{t}(x_{t})\leq h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}(x_{t})$ then the same inequality hold for t1t-1 and the inequality hold for the base case, $t=T$ by definition, hence corollary˜2 is established by induction.
Corollary 2.
The likelihood $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$ is lower bounded by the backward representer $\beta_{t}$:
$h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}(x_{t})\geq\beta_{t}(x_{t}).$
2.3 Reverse-time Markov factorization of the variational family
In this section the factorization (18) is used in combination with forward dynamic programming, where the variational inference problem is formally solved by a forward recursion for a value functional to be determined. Note that the first term of (20) decomposes as
$\mathcal{V}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}t}^{0\mathrel{\mathop{\ordinarycolon}}t},q_{0\mathrel{\mathop{\ordinarycolon}}t})=\int\mathcal{V}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}t}^{0\mathrel{\mathop{\ordinarycolon}}t},q_{0\mathrel{\mathop{\ordinarycolon}}t-1\mid t})q_{t}(x_{t})\operatorname{d\!}x_{t}-\int\log\{q_{t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}$ (27)
The forward value functional is then obtained by maximizing the first term of (27) with respect to $q_{0\mathrel{\mathop{\ordinarycolon}}t-1\mid t}$.
Definition 4 (Variational forward value functional).
Define the variational forward value functional for $0<t\leq T$ by
$V^{\mathsf{f}}_{t}(q_{t})=\max_{q_{0\mathrel{\mathop{\ordinarycolon}}t-1\mid t}\in\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}t-1\mid t}}\int\mathcal{V}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}t}^{0\mathrel{\mathop{\ordinarycolon}}t},q_{0\mathrel{\mathop{\ordinarycolon}}t-1\mid t})q_{t}(x_{t})\operatorname{d\!}x_{t}$ (28)
and define the initial condition, $V^{\mathsf{f}}_{0}$, by
$V^{\mathsf{f}}_{0}(q_{0})=\int\log\{\bar{\pi}_{0}^{0}(x_{0})\}q_{0}(x_{0})\operatorname{d\!}x_{0}.$ (29)
Just as the variational backward value functional satisfies a backward recursion, the variational forward value functional satisfies a forward recursion. Namely, the following identity holds for $1\leq t$
$\mathcal{V}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}t}^{0\mathrel{\mathop{\ordinarycolon}}t},q_{0\mathrel{\mathop{\ordinarycolon}}t-1\mid t})=\int\mathcal{V}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1},q_{0\mathrel{\mathop{\ordinarycolon}}t-2\mid t-1})q_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}+\mathcal{V}(\bar{\pi}_{t\mid t-1}^{t},q_{t-1\mid t})$
where $\mathcal{V}(\bar{\pi}_{0}^{0},q_{0\mathrel{\mathop{\ordinarycolon}}-1\mid 0})=\log\{\bar{\pi}_{0}^{0}(x_{0})\}$ by convention. Inserting this into definition˜4 then gives variational forward value functional at time $t$ according to
$V^{\mathsf{f}}_{t}(q_{t})=\max_{q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}}\Big[V^{\mathsf{f}}_{t-1}(q_{t-1})+\int\mathcal{V}(\bar{\pi}_{t\mid t-1}^{t},q_{t-1\mid t})q_{t}(x_{t})\operatorname{d\!}x_{t}\Big].$
Here $q_{t-1}$ is obtained by marginalizing $q_{t-1\mid t}$ with respect to $q_{t}$. Accounting for this constraint gives the following result.
Lemma 2.
The variational forward value functional satisfies the following recursion
$V^{\mathsf{f}}_{t}(q_{t})=\max_{q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}}\Big[V^{\mathsf{f}}_{t-1}(q_{t-1})+\int\mathcal{V}(\bar{\pi}_{t\mid t-1}^{t},q_{t-1\mid t})q_{t}(x_{t})\operatorname{d\!}x_{t}\Big],$ (30)
where the maximization is subject to the constraint
$q_{t-1}(x_{t-1})=\int q_{t-1\mid t}(x_{t-1}\mid x_{t})q_{t}(x_{t})\operatorname{d\!}x_{t}.$ (31)
Similarly to the variational backward value functional, the variational forward value functional is a linear functional. Namely, make the following ansatz
$V^{\mathsf{f}}_{t}(q_{t})=\int\log\{\alpha_{t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t},\quad 0\leq t\leq T.$ (32)
Then by definition˜4, the initial condition of $\alpha$ is given by $\alpha_{0}=\bar{\pi}_{0}^{0}$. Additionally, if $q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}$ and $q_{t-1}$ satisfies the constraint (31), then the forward value functional is at time $t$ given by
$V^{\mathsf{f}}_{t-1}(q_{t-1})=\int\log\{\alpha_{t-1}(x_{t-1})\}q_{t-1\mid t}(x_{t-1}\mid x_{t})q_{t}(x_{t})\operatorname{d\!}x_{t-1\mathrel{\mathop{\ordinarycolon}}t}$ (33)
Inserting this into the recursion of lemma˜2 gives the following result by induction.
Theorem 4.
The variational forward value functional satisfies the following recursion
$\begin{split}&\int\log\{\alpha_{t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}\\ &=\underset{q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}}{\max}\,\int\Big(\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})\alpha_{t-1}(x_{t-1})}{q_{t-1\mid t}(x_{t-1}\mid x_{t})}\Big\}q_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}\Big)q_{t}(x_{t})\operatorname{d\!}x_{t}.\end{split}$ (34)
Additionally, the terminal density of the variational posterior satisfies
$q_{T}^{0\mathrel{\mathop{\ordinarycolon}}T}=\underset{q_{T}\in\mathcal{Q}_{T}}{\arg\,\max}\,\int\log\Big\{\frac{\alpha_{T}(x_{T})}{q_{T}(x_{T})}\Big\}q_{T}(x_{T})\operatorname{d\!}x_{T},$ (35)
and the lower bound is given by the corresponding maxima.
Theorem˜4 does in principle allow for computing a reverse-time Markov factorization of the variational posterior via a forward recursion. However, just as for the backward recursion the value functional recursion relies on the time-marginals $q_{1},\ldots,q_{T}$, which need to be time-marginals of the variational posterior. Just as for the backward dynamic programming approach a fixed-point iteration can be developed, see proposition˜2. This algorithm can be initialized by a variational filter as discussed in section˜3.
Proposition 2 (Forward variational fixed-point equations).
The forward value functional and the variational posterior satisfy the following fixed-point equations for $1\leq t\leq T$
$\alpha_{0}$ $=\bar{\pi}_{0}^{0}$ (36a)
$q_{t\mid t-1}^{0\mathrel{\mathop{\ordinarycolon}}T}$ $=\underset{q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}}{\arg\,\max}\,\int\Big(\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}\alpha_{t-1}}{q_{t-1\mid t}}\Big\}q_{t-1\mid t}\operatorname{d\!}x_{t-1}\Big)q_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}\operatorname{d\!}x_{t}$ (36b)
$\log\alpha_{t}$ $=\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}\alpha_{t-1}}{q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}}\Big\}q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}\operatorname{d\!}x_{t-1}$ (36c)
$q_{T}^{0\mathrel{\mathop{\ordinarycolon}}T}$ $=\underset{q_{T}\in\mathcal{Q}_{T}}{\arg\,\max}\,\int\log\Big\{\frac{\alpha_{T}}{q_{T}}\Big\}q_{T}\operatorname{d\!}x_{T}$ (36d)
$q_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}T}$ $=\int q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}q_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}\operatorname{d\!}x_{t}.$ (36e)
2.3.1 The forward representer as an unnormalized density
The forward representer, $\alpha_{t}$ is the variational analogues to the unnormalized filtering density, $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$. Similarly, to the backward representer they serve as a lower bound to its analogues. Namely, by Jensen’s inequality
$\begin{split}\log\alpha_{t}(x_{t})&=\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})\alpha_{t-1}(x_{t-1})}{q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{t-1}\mid x_{t})}\Big\}q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}\\ &\leq\log\int\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})\alpha_{t-1}(x_{t-1})\operatorname{d\!}x_{t-1}.\end{split}$ (37)
Therefore, forward representer satisfy the following difference inequality
$\alpha_{t}(x_{t})\leq\int\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})\alpha_{t-1}(x_{t-1})\operatorname{d\!}x_{t-1}.$ (38)
but $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$ satisfy this relation with strict equality. Hence by similar induction argument as for corollary˜2, the following bound can be established.
Corollary 3.
The unnormalized filtering distribution $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$ and the forward value functional satisfy the following bound
$\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})\geq\alpha_{t}(x_{t}),$ (39)
hence, $\alpha_{t}$ is the density of a positive finite measure.
2.3.2 The value functional recursion in predictor-corrector form
It ought to be noted that the forward value functional recursion can be written in so-called predictor-corrector form. Inserting the identity $\bar{\pi}_{t\mid t-1}^{t}=h_{t\mid t}\pi_{t\mid t-1}$ into theorem˜4 gives
$\begin{split}&\int\log\{\alpha_{t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}=\int\log\{h_{t\mid t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}\\ &+\underset{q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}}{\max}\,\int\Big(\int\log\Big\{\frac{\pi_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})\alpha_{t-1}(x_{t-1})}{q_{t-1\mid t}(x_{t-1}\mid x_{t})}\Big\}q_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}\Big)q_{t}(x_{t})\operatorname{d\!}x_{t}.\end{split}$
From which the following result is immediate.
Proposition 3.
The forward value functional satisfy the following predictor-corrector recursion:
$\begin{split}&\int\log\{\alpha_{t}^{\mathsf{p}}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}\\ &=\underset{q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}}{\max}\,\int\Big(\int\log\Big\{\frac{\pi_{t\mid t-1}(x_{t}\mid x_{t-1})\alpha_{t-1}(x_{t-1})}{q_{t-1\mid t}(x_{t-1}\mid x_{t})}\Big\}q_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}\Big)q_{t}(x_{t})\operatorname{d\!}x_{t}\end{split}$
$\log\alpha_{t}(x_{t})=\log\alpha_{t}^{\mathsf{p}}(x_{t})+\log h_{t\mid t}(x_{t}).$
Additionally, a result similar to corollary˜3 for the predictive forward value functional. That is, from Jensen’s inequality and corollary˜3, the following bound hold
$\begin{split}\log\alpha_{t}^{\mathsf{p}}(x_{t})&\leq\log\int\pi_{t\mid t-1}(x_{t}\mid x_{t-1})\alpha_{t-1}(x_{t-1})\operatorname{d\!}x_{t-1}\\ &\leq\log\int\pi_{t\mid t-1}(x_{t}\mid x_{t-1})\bar{\pi}_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}(x_{t-1})\operatorname{d\!}x_{t-1}=\log\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t-1}(x_{t}).\end{split}$ (41)
Corollary 4.
The unnormalized predictive distribution $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t-1}$ and the predictive forward value functional satisfy the following bound
$\alpha_{t}^{\mathsf{p}}(x_{t})\leq\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t-1}(x_{t}).$ (42)
2.4 A Variational two-filter formula
In order to complete the theory of variational state estimation, a relation corresponding to the two-filter formula of corollary˜1 need to be established. By combining (20) and (27), the variational lower bound is obtained as
$\begin{split}\mathcal{V}_{0\mathrel{\mathop{\ordinarycolon}}T}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T},q_{0\mathrel{\mathop{\ordinarycolon}}T})&=\int\mathcal{V}(\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}t}^{0\mathrel{\mathop{\ordinarycolon}}t},q_{0\mathrel{\mathop{\ordinarycolon}}t-1\mid t})q_{t}(x_{t})\operatorname{d\!}x_{t}+\int\mathcal{V}(\bar{\pi}_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}^{t+1\mathrel{\mathop{\ordinarycolon}}T},q_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t})q_{t}(x_{t})\operatorname{d\!}x_{t\mathrel{\mathop{\ordinarycolon}}T}\\ &\quad-\int\log\{q_{t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}.\end{split}$ (43)
Maximizing first with respect to both $q_{0\mathrel{\mathop{\ordinarycolon}}t-1\mid t}$ and $q_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$ and then maximizing with respect to $q_{t}$ gives
$q_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}=\underset{q_{t}\in\mathcal{Q}_{t}}{\arg\,\max}\,\Big[V^{\mathsf{f}}_{t}(q_{t})-\int\log\{q_{t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}+V^{\mathsf{b}}_{t}(q_{t})\Big].$ (44)
Now writing the value functionals in terms of $\alpha_{t}$ and $\beta_{t}$, respectively, gives the variational two-filter formula.
Corollary 5 (Variational two-filter smoother).
The variational smoothing density at time $t$ is the maximizer of
$q_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}=\underset{q_{t}\in\mathcal{Q}_{t}}{\arg\,\max\,}\int\log\Big\{\frac{\alpha_{t}(x_{t})\beta_{t}(x_{t})}{q_{t}(x_{t})}\Big\}q_{t}(x_{t})\operatorname{d\!}x_{t}.$ (45)
3 Optimal and sub-optimal variational filtering
In many applications it is interesting to compute an approximation of the filtering density, $\pi_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$, on-line. Note that the dependence of the value functionals with respect to the observation sequence has hitherto been supressed from the notation. However, in order to appropriately discuss the possibility of variational filtering it needs to be there. That is, let $\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}s}$ denote the forward value functional with respect to the observations $y_{\tau},0\leq\tau\leq s$. Now in view of theorem˜4 and corollary˜3, the variational filtering densities may be defined as follows.
Definition 5 (Optimal variational filtering).
The variational filtering densities, $q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$, with respect to the family $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$, are given by
$q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}=\underset{}{\arg\,\max}\,\int\log\Big\{\frac{\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})}{q_{t}(x_{t})}\Big\}q_{t}(x_{t})\operatorname{d\!}x_{t}.$ (46)
The problem is now immediately clear, theorem˜4 gives a recursion in $t$ of $\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}$. However, in order to obtain a recursive variational filter what is needed is a recursion of $\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$ in $t$. The present analysis suggests that this is in general impossible without re-processing previous observations. This indicates that the optimal filter given by definition˜5 is of time-complexity $O(T^{2})$ (c.f proposition˜2).
However, there is a special case when the optimal variational filter is of time-complexity $O(T)$. Define the universally optimal predictive value functional, $\alpha_{t}^{\dagger}$, and the universally optimal reverse-time transition kernel, $q_{t-1\mid t}^{\dagger}$ with respect to $\alpha_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}$, by Bayes’ rule
$\alpha_{t}^{\dagger}q_{t-1\mid t}^{\dagger}=\pi_{t\mid t-1}\alpha_{t-1}.$ (47)
Inserting this into the recursion of theorem˜4 then gives
$\begin{split}&\int\log\{\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}=\int\log\{h_{t\mid t}(x_{t})\alpha_{t}^{\dagger}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}\\ &+\underset{q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}}{\max}\,\int\Big(\int\log\Big\{\frac{q_{t-1\mid t}^{\dagger}(x_{t-1}\mid x_{t})}{q_{t-1\mid t}(x_{t-1}\mid x_{t})}\Big\}q_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}\Big)q_{t}(x_{t})\operatorname{d\!}x_{t}.\end{split}$
If $q_{t-1\mid t}^{\dagger}\in\mathcal{Q}_{t-1\mid t}$, then the last term on the right-hand side disappears and the following result is evident.
Proposition 4 (Recursive optimal variational filtering).
Assume that $q_{t-1\mid t}^{\dagger}\in\mathcal{Q}_{t-1\mid t}$, then the forward value functional satisfies the following recursion for $1\leq t\leq T$
$\alpha_{t}^{\dagger}(x_{t})$ $=\int\pi_{t\mid t-1}(x_{t}\mid x_{t-1})\alpha_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}(x_{t-1})\operatorname{d\!}x_{t-1}$ (48a)
$\log\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})$ $=\log\{h_{t\mid t}(x_{t})\alpha_{t}^{\dagger}(x_{t})\}.$ (48b)
Additionally, $\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}=\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}$.
Proposition˜4 is of course just a disguised version of the Bayesian filtering recursion of theorem˜2. Therefore, it appears that when $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}t-1\mid t}$ is not rich enough to contain $\pi_{0\mathrel{\mathop{\ordinarycolon}}t-1\mid t}$, then some notion of sub-optimal variational filtering is required. One such candidate is specified in definition˜6.
Definition 6 (Sub-optimal variational filter I).
The variational filter of type I satisfies the following equations for $1\leq t$
$\alpha_{0}^{0}$ $=\bar{\pi}_{0}^{0}$ (49a)
$q_{0}^{0}$ $=\underset{q_{0}\in\mathcal{Q}_{0}}{\arg\,\max}\,\int\log\Big\{\frac{\alpha_{0}^{0}}{q_{0}}\Big\}q_{0}\operatorname{d\!}x_{0}$ (49b)
$q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t},\,q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}$ $=\underset{q_{t}\in\mathcal{Q}_{t},\,q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}}{\arg\,\max}\,\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}\alpha_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}}{q_{t-1\mid t}q_{t}}\Big\}q_{t-1\mid t}q_{t}\operatorname{d\!}x_{t-1\mathrel{\mathop{\ordinarycolon}}t}$ (49c)
$\log\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$ $=\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}\alpha_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}}{q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}}\Big\}q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}\operatorname{d\!}x_{t-1}.$ (49d)
The procedure of definition˜6 is indeed sub-optimal but new definition of $\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$ in fact also provides a lower bound on the unnormalized filtering density $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$ – the argument is exactly the same as for corollary˜3. It is just not necessarily the optimal bound obtainable with respect to the chosen variational family. Furthermore, the sub-optimal variational filter serves as a reasonable method of initializing the fixed-point iteration of proposition˜2.
4 On the intractability of the value functionals
The fact that the value functional recursions of theorem˜1, theorem˜2, and definition˜6 are in general intractable has hitherto not been discussed. However, $\alpha_{t}$ ought to be an unnormalized density (see corollary˜3). Therefore, some care needs to taken to ensure that $\alpha_{t}$ remains normalizable throughout recursion. Here this issue shall not be examined in exhaustive detail, there is simply too many possible approaches that will have to be explored elsewhere. However, there is a simple approach approach will be offered. Namely, by definition˜6, the normalization constant of $\alpha_{t}$ has a lower bound according to
$\begin{split}\log\int\alpha_{t}(x_{t})\operatorname{d\!}x_{t}&\geq\int\log\Big\{\frac{\alpha_{t}(x_{t})}{q_{t}(x_{t})}\Big\}q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})\operatorname{d\!}x_{t}\\ &=\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})\alpha_{t-1}(x_{t-1})}{q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{t-1}\mid x_{t})q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})}\Big\}q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}(x_{t-1}\mid x_{t})q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})\operatorname{d\!}x_{t-1\mathrel{\mathop{\ordinarycolon}}t}.\end{split}$ (50)
This then serves as an approximation of the log-normalizing constant of $\alpha_{t}$, combining this with the fact that $q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$ is the best approximation of the normalization of $\alpha_{t}$ gives the following approximation.
Approximation 1 (Variational approximation of the forward value functional I).
The forward value functional is approximately given by
$\log\hat{\kappa}_{t}$ $=\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}\alpha_{t-1}}{q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})}\Big\}q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})\operatorname{d\!}x_{t-1\mathrel{\mathop{\ordinarycolon}}t}$ (51a)
$\log\alpha_{t}$ $\approx\log\hat{\kappa}_{t}+\log q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t}).$ (51b)
The filter proposed by Courts2021 is obtained when the sub-optimal variational filter of definition˜6 is combined with ˜1. Namely, by plugging in ˜1 into definition˜6, the maximization problem of the filter update reads
$\underset{q_{t-1}\in\mathcal{Q}_{t-1},\,q_{t\mid t-1}\in\mathcal{Q}_{t\mid t-1}}{\max}\,\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}q_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}}{q_{t\mid t-1}q_{t-1}}\Big\}q_{t-1\mid t}q_{t}\operatorname{d\!}x_{t-1\mathrel{\mathop{\ordinarycolon}}t},$ (52)
where Bayes’ rule was used, $q_{t-1\mid t}q_{t}=q_{t\mid t-1}q_{t-1}$. This is precisely algorithm 2 of Courts2021.
5 Related work
Recursive formulations have previously been studied by primarily Weber2015 and Campbell2021 from the reinforcement learning perspective. These recursions are developed in terms of certain value functions that of course are related to the present notion of value functionals. More precisely, Weber2015 defines a backward value function by
$\tilde{V^{\mathsf{b}}}_{t}(x_{t})=\int\log\Big\{\frac{\bar{\pi}_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}^{t+1\mathrel{\mathop{\ordinarycolon}}T}(x_{t+1\mathrel{\mathop{\ordinarycolon}}T}\mid x_{t})}{q_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}(x_{t+1\mathrel{\mathop{\ordinarycolon}}T}\mid x_{t})}\Big\}q_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}(x_{t+1\mathrel{\mathop{\ordinarycolon}}T}\mid x_{t})\operatorname{d\!}x_{t+1\mathrel{\mathop{\ordinarycolon}}T}.$
The following relation is then formally identified (see definition˜3)
$\tilde{V^{\mathsf{b}}}_{t}(x_{t})=\log\beta_{t}(x_{t}).$
However, the relation between $\beta_{t}$ and $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$ provided by corollary˜2 was omitted by Weber2015 and so was the fixed-point equations of proposition˜2. On the other hand, Campbell2021 defines a forward value function by
$\tilde{V^{\mathsf{f}}}_{t}(x_{t})=\int\log\Big\{\frac{\bar{\pi}_{0\mathrel{\mathop{\ordinarycolon}}t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{0\mathrel{\mathop{\ordinarycolon}}t})}{q_{0\mathrel{\mathop{\ordinarycolon}}t}(x_{0\mathrel{\mathop{\ordinarycolon}}t})}\Big\}q_{0\mathrel{\mathop{\ordinarycolon}}t-1\mid t}(x_{0\mathrel{\mathop{\ordinarycolon}}t-1}\mid x_{t})\operatorname{d\!}x_{0\mathrel{\mathop{\ordinarycolon}}t-1}.$
In view of definition˜4 and theorem˜4 this value function may formally be identified as
$\tilde{V^{\mathsf{f}}}_{t}(x_{t})=\log\Big\{\frac{\alpha_{t}(x_{t})}{q_{t}(x_{t})}\Big\}.$
However, the relation between $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$ and $\alpha_{t}$ established by corollary˜3 was not provided by Campbell2021, nor was the fixed-point equations proposition˜2. Furthermore, they put forth the following value function recursion
$\begin{split}\tilde{V^{\mathsf{f}}}_{t+1}(x_{t+1})&=\int\log\Big\{\frac{\bar{\pi}_{t+1\mid t}^{t+1}(x_{t+1}\mid x_{t})q_{t}(x_{t})}{q_{t\mid t+1}(x_{t}\mid x_{t+1})q_{t+1}(x_{t+1})}\Big\}q_{t\mid t+1}(x_{t}\mid x_{t+1})\operatorname{d\!}x_{t}\\ &\quad+\int\tilde{V^{\mathsf{f}}}_{t}(x_{t})q_{t\mid t+1}(x_{t}\mid x_{t+1})\operatorname{d\!}x_{t}.\end{split}$ (53)
Now if $q_{t}$ actually satisfies the following constraint
$q_{t}(x_{t})=\int q_{t\mid t+1}(x_{t}\mid x_{t+1})q_{t+1}(x_{t+1}),$
then the value function recursion of (53) is formally equivalent to theorem˜5, namely
$\tilde{V^{\mathsf{f}}}_{t+1}(x_{t+1})=\int\log\Big\{\frac{\bar{\pi}_{t+1\mid t}^{t+1}(x_{t+1}\mid x_{t})\alpha_{t}(x_{t})}{q_{t\mid t+1}(x_{t}\mid x_{t+1})q_{t+1}(x_{t+1})}\Big\}q_{t\mid t+1}(x_{t}\mid x_{t+1})\operatorname{d\!}x_{t}.$
However, this is not the approach suggested by Campbell2021. Instead they identify $q_{t}$ with their corresponding filtering distribution $q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$. Notably, the present approach circumvents this issue by recursing over $\alpha_{t}$ rather than $\tilde{V^{\mathsf{f}}}_{t}$. This also means that the present notion of variational filtering is different from that of Campbell2021.
A few days later after the first pre-print of the present manuscript was released, another approach appeared by (Abdulsamad2025). Where they study iterative recursive variational Gaussian smoothing under entropic regularization using a Lagrange multiplier approach rather than dynamic programming. Also, rather than imposing a constraint class $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ explicitly they “close” the value functional recursions by a variety of quadratic approximations.
6 Case study: Jump Gauss–Markov systems
In this section the variational forward method is studied in jump Gauss–Markov systems under a certain factored approximation, which leads to an alternate coordinate ascent method for computing the transition kernels. More specifically, consider the following Markov chain
$z_{0}$ $\sim\lambda_{0}(z_{0})$ (54a)
$z_{t}\mid z_{t-1}$ $\sim\lambda_{t\mid t-1}(z_{t}\mid z_{t-1}),$ (54b)
where $z_{t}$ takes values in $\{1,\ldots,M\}$. The process $z_{t}$ drives two other processes $x_{t}$ and $y_{t}$ according to
$x_{0}\mid z_{0}$ $\sim\rho_{0}(x_{0}\mid z_{0})$ (55a)
$x_{t}\mid x_{t-1}\,,z_{t-1}$ $\sim\rho_{t\mid t-1}(x_{t}\mid x_{t-1},z_{t-1})$ (55b)
$y_{t}\mid x_{t}\,,z_{t}$ $\sim h_{t\mid t}(y_{t}\mid x_{t},z_{t}),$ (55c)
Additionally, a conditionally Gaussian assumption is imposed on $x_{t}$ and $y_{t}$, namely
$\rho_{0}(x_{0}\mid z_{0})$ $=\mathcal{N}(x_{0};\mu_{0}(z_{0}),\Sigma_{0}(z_{0}))$ (56a)
$a_{t\mid t-1}(x_{t-1},z_{t-1})$ $=\Phi_{t,t-1}(z_{t-1})x_{t-1}+u_{t,t-1}(z_{t-1})$ (56b)
$\rho_{t\mid t-1}(x_{t}\mid x_{t-1},z_{t-1})$ $=\mathcal{N}(x_{t};a_{t\mid t-1}(x_{t-1},z_{t-1}),Q_{t\mid t-1}(z_{t-1}))$ (56c)
$c_{t\mid t}(x_{t},z_{t})$ $=C_{t\mid t}(z_{t})x_{t}+b_{t\mid t}(z_{t})$ (56d)
$h_{t\mid t}(y_{t}\mid x_{t},z_{t})$ $=\mathcal{N}(y_{t};c_{t\mid t}(x_{t},z_{t}),R_{t\mid t}(z_{t})).$ (56e)
Taken together, the triplet $(z_{t},x_{t},y_{t})$ forms a jump Gauss–Markov system and the goal here is to infer the complete hidden state $(z_{t},x_{t})$ given the observed sequence $y_{0\mathrel{\mathop{\ordinarycolon}}T}$. The transition kernel of the complete hidden state is then
$\pi_{t\mid t-1}(x_{t},z_{t}\mid x_{t-1},z_{t-1})=\rho_{t\mid t-1}(x_{t}\mid x_{t-1},z_{t-1})\lambda_{t\mid t-1}(z_{t}\mid z_{t-1}).$ (57)
In this section the variational state estimation methodology is applied under a factorized variational approximation according to
$q_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T},z_{0\mathrel{\mathop{\ordinarycolon}}T})=f_{0\mathrel{\mathop{\ordinarycolon}}T}(z_{0\mathrel{\mathop{\ordinarycolon}}T})g_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T}).$ (58)
The aim of this section is to show, under the factorized class (58), that the forward value functional satisfy the following theorem.
Theorem 5 (Tractable forward value functional).
The forward value functional is for $0\leq t\leq T$ respectively given by
$\log\alpha_{t}(x_{t},z_{t})=\log\kappa_{t}+\log f^{\star}_{t}(z_{t})+\log g^{\star}_{t}(x_{t}\mid z_{t}),$ (59)
where $g^{\star}_{t}$ is given by
$g^{\star}_{t}(x_{t}\mid z_{t})=\mathcal{N}(x_{t};\mu^{\star}_{t}(z_{t}),\Sigma^{\star}_{t}(z_{t})).$ (60)
Additionally, the optimal $g_{0\mathrel{\mathop{\ordinarycolon}}T}$ is Gauss–Markov.
Demonstrating the veracity of theorem˜5 and developing the value functional recurrences will be done in several steps. Before doing so, the filtering update will be established. Namely, let $\xi_{t}$ be the negative relative entropy between $g_{t}$ and $g_{t}^{\star}$, that is
$\begin{split}\log\xi_{t}(z_{t})&=\int\log\Big\{\frac{g^{\star}_{t}(x_{t}\mid z_{t})}{g_{t}(x_{t})}\Big\}g_{t}(x_{t})\operatorname{d\!}x_{t}.\end{split}$ (61)
Then in view of definition˜6, the filtering distributions $g_{t}$ and $f_{t}$ maximize:
$\sum_{z_{t}}\Big(\int\log\Big\{\frac{\alpha_{t}(x_{t},z_{t})}{g_{t}(x_{t})f_{t}(z_{t})}\Big\}g_{t}(x_{t})\operatorname{d\!}x_{t}\Big)f_{t}(z_{t})=\sum_{z_{t}}\log\Big\{\frac{\xi_{t}(z_{t})f_{t}^{\star}(z_{t})}{f_{t}(z_{t})}\Big\}f_{t}(z_{t})+\log\kappa_{t}.$
Clearly, the stationary point with respect to $f_{t}$ is given by $f_{t}\propto\xi_{t}f_{t}^{\star}$. Similarly, the stationary point with respect to $g_{t}$ satisfy
$g_{t}(x_{t})\propto\exp\Big(\sum_{z_{t}}\log\{\alpha_{t}(x_{t},z_{t})\}f_{t}(z_{t})\Big)$
and by theorem˜5 $g_{t}$ is then log-quadratic in $x_{t}$. That is to say Gaussian, the parameters of which are obtained by simple algebra. The following theorem is obtained.
Theorem 6.
If theorem˜5 hold, then the filtering distributions $g_{t}$ and $f_{t}$ satisfy the following fixed-point equations
$f_{t}(z_{t})$ $=\frac{\xi_{t}(z_{t})f_{t}^{\star}(z_{t})}{\sum_{z}\xi_{t}(z)f_{t}^{\star}(z)}$ (62a)
$\Sigma_{t}^{-1}$ $=\sum_{z_{t}}\{\Sigma^{\star}_{t}(z_{t})\}^{-1}f_{t}(z_{t})$ (62b)
$\Sigma_{t}^{-1}\mu_{t}$ $=\sum_{z_{t}}\{\Sigma^{\star}_{t}(z_{t})\}^{-1}\mu^{\star}_{t}(z_{t})f_{t}(z_{t}).$ (62c)
Additionally, the marginal likelihood is lower bounded by
$\log h_{0\mathrel{\mathop{\ordinarycolon}}t}\geq\log\Big\{\sum_{z}\xi_{t}(z)f_{t}^{\star}(z)\Big\}+\log\kappa_{t}.$ (63)
6.1 The variational forward method
For the induction step, assume that (59) holds at t1t-1 in the following. Start by applying Bayes’ rule to $\lambda_{t\mid t-1}$ and $f^{\star}_{t-1}$, which gives the following identity
$\lambda_{t\mid t-1}(z_{t}\mid z_{t-1})f^{\star}_{t-1}(z_{t-1})=f_{t}^{\mathsf{p}}(z_{t})f_{t-1\mid t}^{\dagger}(z_{t-1}\mid z_{t})$ (64)
Similarly, the following identity also hold by Bayes’ rule
$\rho_{t\mid t-1}(x_{t}\mid x_{t-1},z_{t-1})g^{\star}_{t-1}(x_{t-1}\mid z_{t-1})$ $=\tilde{g}^{\mathsf{p}}_{t}(x_{t}\mid z_{t-1})\tilde{g}_{t-1\mid t}(x_{t-1}\mid x_{t},z_{t-1})$ (65a)
$\tilde{g}^{\mathsf{p}}_{t}(x_{t}\mid z_{t-1})$ $=\mathcal{N}(x_{t};\tilde{\mu}^{\mathsf{p}}_{t}(z_{t-1}),\tilde{\Sigma}^{\mathsf{p}}_{t}(z_{t-1}))$ (65b)
$\tilde{g}_{t-1\mid t}(x_{t-1}\mid x_{t},z_{t-1})$ $=\mathcal{N}(x_{t-1};\tilde{a}_{t-1\mid t}(x_{t},z_{t-1}),\tilde{Q}_{t-1\mid t}(z_{t-1})),$ (65c)
where $\tilde{g}^{\mathsf{p}}_{t}$ , and $\tilde{g}_{t-1\mid t}$ are obtained by Gaussian marginalization and conditioning, respectively. Inserting this into theorem˜4 gives the following result.
Lemma 3.
If (59) hold at t1t-1, then the forward value functional at time $t$ is given by
$\begin{split}&\log\alpha_{t}(x_{t},z_{t})=\log f_{t}^{\mathsf{p}}(z_{t})+\sum_{z_{t-1}}\log\Big\{\frac{f_{t-1\mid t}^{\dagger}(z_{t-1}\mid z_{t})}{f_{t-1\mid t}(z_{t-1}\mid z_{t})}\Big\}f_{t-1\mid t}(z_{t-1}\mid z_{t})+\log h_{t\mid t}(x_{t},z_{t})\\ &+\sum_{z_{t-1}}\Big(\int\log\Big\{\frac{\tilde{g}_{t-1\mid t}(x_{t-1}\mid x_{t},z_{t-1})}{g_{t-1\mid t}(x_{t-1}\mid x_{t})}\Big\}g_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}\Big)f_{t-1\mid t}(z_{t-1}\mid z_{t})\\ &+\log\kappa_{t-1}+\sum_{z_{t-1}}\log\{\tilde{g}^{\mathsf{p}}_{t}(x_{t}\mid z_{t-1})\}f_{t-1\mid t}(z_{t-1}\mid z_{t}).\end{split}$
Lemma˜3 will be useful for obtaining a fixed-point equation for the reverse-time transition kernels in the sequel. Before doing so, define the likelihood $\eta_{t}$, the Gaussian transiton kernel $g^{\dagger}_{t-1\mid t}$, and the Gaussian density $g^{\mathsf{p}}_{t}$ by  [^3] A convenient way of organizing of $\eta_{t}$, $g^{\dagger}_{t-1\mid t}$, and $g^{\mathsf{p}}_{t}$ computation is provided by lemma 5 in appendix A
$\begin{split}&\log\{\eta_{t}(z_{t})g^{\dagger}_{t-1\mid t}(x_{t-1}\mid x_{t},z_{t})g^{\mathsf{p}}_{t}(x_{t}\mid z_{t})\}\\ &\quad=\sum_{z_{t-1}}\log\{\tilde{g}^{\mathsf{p}}_{t}(x_{t}\mid z_{t-1})\tilde{g}_{t-1\mid t}(x_{t-1}\mid x_{t},z_{t-1})\}f_{t-1\mid t}(z_{t-1}\mid z_{t}).\end{split}$ (66)
Additionally, define the likelihoods $\zeta_{t}^{\dagger}$, $\eta_{t}^{\dagger}$ and $h_{t\mid t}^{\dagger}$ by  [^4] A convenient way of organizing the computation of $\eta_{t}^{\dagger}$ and $h_{t\mid t}^{\dagger}$ is provided by lemma 6 in appendix A.
$\log\zeta_{t}^{\dagger}(z_{t})$ $=\sum_{z_{t-1}}\log\Big\{\frac{f_{t-1\mid t}^{\dagger}(z_{t-1}\mid z_{t})}{f_{t-1\mid t}(z_{t-1}\mid z_{t})}\Big\}f_{t-1\mid t}(z_{t-1}\mid z_{t})$ (67a)
$\log\{\eta_{t}^{\dagger}(z_{t})h_{t\mid t}^{\dagger}(x_{t},z_{t})\}$ $=\int\log\Big\{\frac{g^{\dagger}_{t-1\mid t}(x_{t-1}\mid x_{t},z_{t})}{g_{t-1\mid t}(x_{t-1}\mid x_{t})}\Big\}g_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1},$ (67b)
where it can be seen that $h_{t\mid t}^{\dagger}$ is indeed a Gaussian likelihood. Then by successive substitutions, the forward value functional is given by
$\begin{split}\log\alpha_{t}(x_{t},z_{t})&=\log\{\eta_{t}^{\dagger}(z_{t})\zeta_{t}^{\dagger}(z_{t})\eta_{t}(z_{t})f_{t}^{\mathsf{p}}(z_{t})\}\\ &\quad+\log\{h_{t\mid t}^{\dagger}(x_{t},z_{t})h_{t\mid t}(x_{t},z_{t})g^{\mathsf{p}}_{t}(x_{t}\mid z_{t})\}+\log\kappa_{t-1}.\end{split}$
The forward value functional update is now readily obtained and the proof of theorem˜5 is complete.
Proposition 5 (Forward value functional update).
Given that $\alpha_{t-1}$ is of the form stated in theorem˜5, then so is $\alpha_{t}$. The update is given by the following relations:
$\bar{h}_{t\mid t}(z_{t})g^{\star}_{t}(x_{t}\mid z_{t})$ $=h_{t\mid t}^{\dagger}(x_{t},z_{t})h_{t\mid t}(x_{t},z_{t})g^{\mathsf{p}}_{t}(x_{t}\mid z_{t})$ (68a)
$\kappa_{t\mid t-1}$ $=\sum_{z_{t}}\bar{h}_{t\mid t}(z_{t})\eta_{t}^{\dagger}(z_{t})\zeta_{t}^{\dagger}(z_{t})\tilde{\eta}_{t}(z_{t})f_{t}^{\mathsf{p}}(z_{t})$ (68b)
$f_{t}^{\star}(z_{t})$ $=\kappa_{t\mid t-1}^{-1}\bar{h}_{t\mid t}(z_{t})\eta_{t}^{\dagger}(z_{t})\zeta_{t}^{\dagger}(z_{t})\tilde{\eta}_{t}(z_{t})f_{t}^{\mathsf{p}}(z_{t})$ (68c)
$\log\kappa_{t}$ $=\log\kappa_{t-1}+\log\kappa_{t\mid t-1},$ (68d)
where the first relation is Bayes’ rule.
6.1.1 Fixed-point equations for the transition kernels
It remains to obtain the fixed-point equations for the reverse-time transition kernels $g_{t-1\mid t}$ and $f_{t-1\mid t}$. They are given by as the maximizers of
$\sum_{z_{t}}\Big(\int\log\alpha_{t}(x_{t},z_{t})g_{t}(x_{t})\operatorname{d\!}x_{t}\Big)f_{t}(z_{t}).$
In view of lemma˜3, define $\zeta_{t-1}$ by
$\begin{split}\log\zeta_{t-1}(z_{t-1})&=\int\Big(\int\log\Big\{\frac{\tilde{g}_{t-1\mid t}(x_{t-1}\mid x_{t},z_{t-1})}{g_{t-1\mid t}(x_{t-1}\mid x_{t})}\Big\}g_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}\Big)g_{t}(x_{t})\operatorname{d\!}x_{t}\\ &\quad+\int\log\{\tilde{g}^{\mathsf{p}}_{t}(x_{t}\mid z_{t-1})\}g_{t}(x_{t})\operatorname{d\!}x_{t},\end{split}$ (69)
then it is immediately clear that $f_{t-1\mid t}\propto\zeta_{t-1}f_{t-1\mid t}^{\dagger}$. Additionally, by lemma˜3 it is clear that $g_{t-1\mid t}$ is given as the maximizer of
$\sum_{z_{t},z_{t-1}}\Big(\int\log\Big\{\frac{\tilde{g}_{t-1\mid t}(x_{t-1}\mid x_{t},z_{t-1})}{g_{t-1\mid t}(x_{t-1}\mid x_{t})}\Big\}g_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}\Big)f_{t-1\mid t}(z_{t-1}\mid z_{t})f_{t}(z_{t}).$
Consequently, the reverse-time transition kernel $g_{t-1\mid t}$ is clearly given by
$g_{t-1\mid t}(x_{t-1}\mid x_{t})\propto\exp\Big(\sum_{z_{t},z_{t-1}}\log\{\tilde{g}_{t-1\mid t}(x_{t-1}\mid x_{t},z_{t-1})\}f_{t-1\mid t}(z_{t-1}\mid z_{t})f_{t}(z_{t})\Big)$ (70)
and since $\tilde{g}_{t-1\mid t}$ is log-quadratic in $x_{t-1}$ and $x_{t}$ then so is $g_{t-1\mid t}$, Gaussian that is.
Proposition 6 (Fixed-point equations fo reverse-time transition kernels).
The reverse-time transiton kernels satisfy the following fixed-point equations
$f_{t-1\mid t}(z_{t-1}\mid z_{t})$ $=\frac{\zeta_{t-1}(z_{t-1})f_{t-1\mid t}^{\dagger}(z_{t-1}\mid z_{t})}{\sum_{z_{t-1}}\zeta_{t-1}(z_{t-1})f_{t-1\mid t}^{\dagger}(z_{t-1}\mid z_{t})}$ (71)
$Q_{t-1\mid t}^{-1}$ $=\sum_{z_{t-1},z_{t}}\tilde{Q}_{t-1\mid t}^{-1}(z_{t-1})f_{t-1\mid t}(z_{t-1}\mid z_{t})f_{t}(z_{t})$ (72)
$Q_{t-1\mid t}^{-1}a_{t-1\mid t}(x)$ $=\sum_{z_{t-1},z_{t}}\tilde{Q}_{t-1\mid t}^{-1}(z_{t-1})a_{t-1\mid t}(x,z_{t-1})f_{t-1\mid t}(z_{t-1}\mid z_{t})f_{t}(z_{t})$ (73)
$g_{t-1\mid t}(x_{t-1}\mid x_{t})$ $=\mathcal{N}(a_{t-1\mid t}(x),Q_{t-1\mid t}).$ (74)
6.2 Simulation study: an autoregressive staircase
Figure 1: One realization of the jump Gauss–Markov system. The markers for $x_{t}$ are colored according to the corresponding value of $z_{t}$.
The model index, $z_{t}$, starts at the uniform distribution on $\{1,\ldots,M\}$ and then stays in its current state with probability $p$ or jumps to any adjacent state with equal probability. The Markov chain is then specified by the following initial and transition probabilities,
$\lambda_{0}(z)$ $=\frac{1}{M}$ (75a)
$\lambda_{t\mid t-1}(z\mid 1)$ $=p\delta_{z,1}+(1-p)\delta_{z,2}$ (75b)
$\lambda_{t\mid t-1}(z\mid z^{\prime})$ $=p\delta_{z,z^{\prime}}+\frac{1-p}{2}(\delta_{z-1,z^{\prime}}+\delta_{z+1,z^{\prime}}),\quad z^{\prime}=2,\ldots,M-1,$ (75c)
$\lambda_{t\mid t-1}(z\mid M)$ $=p\delta_{z,M}+(1-p)\delta_{z,M-1}$ (75d)
where $\delta$ denotes Kronecker’s delta. The state variable switches between $M$ different first order, stationary autoregressive processes with parameters given by
$\mu_{0}(z)$ $=z-1$ (76a)
$Q_{t\mid t-1}$ $=(1-\mathinner{\!\left\lvert\phi_{0}\right\rvert}^{2})\Sigma_{0}(z)$ (76b)
$a_{t\mid t-1}(x,z)$ $=\phi_{0}x+(1-\phi_{0})\mu_{0}(z),$ (76c)
where $\mu_{0}(z)$ and $\Sigma_{0}(z)$ are the stationary means and variances, repectively. Lastly, the parameters of the observation model are given by
$c_{t\mid t}(x,z)$ $=x$ (77a)
$R_{t\mid t}(x,z)$ $=R.$ (77b)
There are various recursive approaches to computing filtering densities and smoothing marginals in jump Gauss–Markov systems (BarShalom2004, Helmick1995, Balenzuela2022). However, to the authors knowledge, none that can recursively compute a posterior over the entire path $(x_{0\mathrel{\mathop{\ordinarycolon}}T},z_{0\mathrel{\mathop{\ordinarycolon}}T})$. In any case, the posterior obtained from proposition˜2 is optimal in relative entropy sense with respect to the chosen variational family. Thus instead the sub-optimal variational filter (VJGM(0)) is compared to an interacting multiple model filter with $M$ components (IMM) (BarShalom2004), to ensure the notion of sub-optimal vairational filtering is reasonable. Additionally, (VJGM(0)) is used to initialize the fixed-point iteration of proposition˜2, which runs for 10 iterations (VJGM(10)).  [^5] It seems to always converge in less than 10 iterations in this example. The posteriors obtained from (VJGM(0)) and (VJGM(10)) are then compared. The parameters are set to $\mu_{0}(z)=z-1$, $\Sigma_{0}(z)=0.5^{2}$, $\phi_{0}=0.5$, $M=4$, $p=0.9$, $R=1$, and the number of time points is set to $T=2^{9}+1$. A realization of the system is shown in fig.˜1.
Figure 2: Performance comparison of the filtering distributions produced by VJGM(0) and IMM. For the $\chi^{2}$-statistic, the corresponding ground-truth density is shown. It can be seen that their performance is quite similar across metrics except for the $\chi^{2}$-square statistic, where IMM significantly outperforms VJGM(0).
The comparisons are carried out over 1000 Monte Carlo trials where the root mean-square errors in estimating $x_{t}$ are recorded along with the mean accuracy and mean log-odds of the correct $z_{t}$ under the filtering distribution when comparing VJGM(0) with IMM and under the smoothing distribution when comparing VJGM(0) with VJGM(10). Additionally, in order to assess the calibration quality in the estimates of $x_{t}$ the sum of $\chi^{2}$-statistics is recorded when comparing VJGM(0) with IMM and the $\chi^{2}$-statistic of the entire sequence $x_{0\mathrel{\mathop{\ordinarycolon}}T}$ when comparing VJGM(0) and VJGM(10).  [^6] When the Gaussian assumption hold both $\chi^{2}$ metrics are realizations from $\chi^{2}(T+1)$.
The results of the comparison between VJGM(0) and IMM are shown in fig.˜2. It can be seen that their performance is similar across all metrics except the $\chi^{2}$-statistic, where the uncertainty in $x_{t}$ produced by VJGM(0) is very overconfident. This might be due to the well documented mode seeking behavior of variational inference (Bishop2006).
Figure 3: Performance comparison of posterior distributions over the path $(x_{0\mathrel{\mathop{\ordinarycolon}}T},z_{0\mathrel{\mathop{\ordinarycolon}}T})$ produced by VJGM(0) and VJGM(10). For the $\chi^{2}$-statistic, the corresponding ground-truth density is shown. It can be seen that VJGM(10) offers a moderate improvement over VJGM(0) in terms of estimation accuracy, calibration, and classification accuracy.
The result of the comparison between VJGM(0) and VJGM(10) are shown in fig.˜3. Here it can be seen that while VJGM(0) produces a posterior of reasonable quality, actually carrying out the fixed-point iteration of proposition˜2 can bring signifcant improvements. Comparing to the filtering results of fig.˜2, it can also be seem that the expected improvement in estimation performance between smoothing and filtering has been realized, particularly in terms of estimating the model index, $z_{t}$.
7 Case study: log-polynomial systems
In this section, variational filtering is studied when the problem is log-polynomial of order, say $\nu$. The model is then given by
$\log\bar{\pi}_{0}^{0}(x_{0})$ $=\log\mathcal{N}(x_{0};\mu_{0},\Sigma_{0})+\sum_{m=0}^{\nu}\ell_{0}^{(m)}x_{0}^{m}$ (78a)
$\log\bar{\pi}_{t\mid t-1}^{t}(x_{t}\mid x_{t-1})$ $=\log\mathcal{N}(x_{t};a_{t\mid t-1}(x_{t-1}),Q_{t\mid t-1})+\sum_{m=0}^{\nu}\ell_{t}^{(m)}x_{t}^{m}$ (78b)
and the state is thus a scalar. Note that by defining $p_{m}(x)=x^{m}$ the likelihoods can be written as inner products according to
$\log h_{t\mid t}(x_{t})=\sum_{m=0}^{\nu}\ell_{t}^{(m)}x_{t}^{m}=\langle\ell_{t},p(x_{t})\rangle.$ (79)
The aim of this section is to establish the forward value functional recursion when $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ is Gauss–Markov and thus the following theorem.
Theorem 7.
Let the variational family $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ be Gauss–Markov then the forward value functional associated with a log-polynomial system of order $\nu\geq 2$ is given by
$\log\alpha_{t}(x_{t})=\sum_{m=0}^{\nu}\lambda_{t}^{(m)}x_{t}^{m}=\langle\lambda_{t},p(x_{t})\rangle.$ (80)
Theorem˜7 readily gives the filter update, namely
$\begin{split}\int\log\Big\{\frac{\alpha_{t}(x_{t})}{q_{t}(x_{t})}\Big\}q_{t}(x_{t})\operatorname{d\!}x_{t}&=\int\langle\lambda_{t},p(x_{t})\rangle q_{t}(x_{t})\operatorname{d\!}x_{t}-\int\log\{q_{t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}.\end{split}$ (81)
Applying lemma˜8 from appendix˜B to the first term then gives the following result.
Proposition 7 (Variational filter update).
Let $q_{t}$ ge Gaussian with mean and covariance $\mu_{t}$ and $\Sigma_{t}$, respectively, then
$\begin{split}\int\log\Big\{\frac{\alpha_{t}(x_{t})}{q_{t}(x_{t})}\Big\}q_{t}(x_{t})\operatorname{d\!}x_{t}&=\langle\lambda_{t},A_{1}(\Sigma_{t})p(\mu_{t})\rangle+\frac{1}{2}(\log\{2\pi\Sigma_{t}\}+1),\end{split}$ (82)
where $A_{1}$ is defined in lemma˜8. Hence the variational filter mean and covariance satisfies
$\mu_{t}^{0\mathrel{\mathop{\ordinarycolon}}t},\,\Sigma_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}=\underset{\mu_{t},\Sigma_{t}}{\arg\,\max}\,\big\{\langle\lambda_{t},A_{1}(\Sigma_{t})p(\mu_{t})\rangle+\frac{1}{2}(\log\{2\pi\Sigma_{t}\}+1)\big\}.$ (83)
It now remains to establish theorem˜7 and a recursion for the value functional. Note that theorem˜7 hold at $t=0$ by the following identification
$\lambda_{0}^{(0)}$ $=-\frac{1}{2}\log\{2\pi\Sigma_{0}\}-\frac{\mu_{0}^{2}}{2\Sigma_{0}}+\ell_{0}^{(0)}$ (84a)
$\lambda_{0}^{(1)}$ $=\frac{\mu_{0}}{\Sigma_{0}}+\ell_{0}^{(1)}$ (84b)
$\lambda_{0}^{(2)}$ $=-\frac{1}{2\Sigma_{0}}+\ell_{0}^{(2)}$ (84c)
$\lambda_{0}^{(m)}$ $=\ell_{0}^{(m)},\quad 2\leq m.$ (84d)
It then remains to establish the induction step. The value functional recursion is
[…]
where (79) was used for the first term and $c_{t}$ is defiend by lemma˜9 in appendix˜B. Additionally, by lemma˜8 in appendix˜B, there are matrices $A_{1}(Q_{t-1\mid t})$ and $A_{2}(\Phi_{t-1,t},u_{t-1,t})$ such that
$\begin{split}\int\langle\lambda_{t-1},p(x_{t-1})\rangle q_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}&=\langle\lambda_{t-1},A_{1}(Q_{t-1\mid t})p(a_{t-1\mid t}(x_{t}))\rangle\\ &=\langle\lambda_{t-1},A_{1}(Q_{t-1\mid t})A_{2}(\Phi_{t-1,t},u_{t-1,t})p(x_{t})\rangle\end{split}$ (86)
Theorem˜7 has then been established along with the following value functional recursion.
Proposition 8 (Forward value functional update).
Assume theorem˜7 hold at t1t-1 then it also hold at $t$. Additionally the parameter $\lambda_{t}$ can be obtained from $\lambda_{t-1}$ according to
$\lambda_{t}=A_{2}^{*}(\Phi_{t-1,t},u_{t-1,t})A_{1}^{*}(Q_{t-1\mid t})\lambda_{t-1}+c_{t}(\Phi_{t-1,t},u_{t-1,t},Q_{t-1\mid t})+\ell_{t}$ (87)
where the matrices $A_{1}$ and $A_{2}$ are defined in lemma˜8, and $c_{t}$ is defined in lemma˜9.
7.1 Simulation study: The cubic sensor problem
Let $a_{t\mid t-1}(x)=\phi_{0}x+(1-\phi_{0})\mu_{0}$ and $c_{t\mid t}(x)=\beta x^{3}$ and consider the cubic sensor system given by (Katayama2013)
$x_{0}$ $\sim\mathcal{N}(\mu_{0},\Sigma_{0})$ (88a)
$x_{t}\mid x_{t-1}$ $\sim\mathcal{N}(x_{t};a_{t\mid t-1}(x_{t-1}),Q_{t\mid t-1})$ (88b)
$y_{t}\mid x_{t}$ $\sim\mathcal{N}(y_{t};c_{t\mid t}(x_{t}),R).$ (88c)
It is clearly a log-polynomial system of order $6$ with the likelihoods given by
$\begin{split}\log h_{t\mid t}(x_{t})&=-\frac{1}{2}\Big(\log\{2\pi R\}+\frac{y_{t}^{2}}{2R}\Big)+\frac{y_{t}\beta}{R}x_{t}^{3}-\frac{\beta^{2}}{2R}x_{t}^{6}\\ &=\sum_{m=0}^{6}\ell_{t}^{(m)}x_{t}^{m}.\end{split}$ (89)
It is interesting to note that this sytsem has two regimes. Namely, $(i)$ $\mathinner{\!\left\lvert x_{t}\right\rvert}<1$ where the signal-to-noise ratio is small and $(ii)$ $\mathinner{\!\left\lvert x_{t}\right\rvert}\geq 1$ where the signal-to-noise-ratio is high. In particular, if $\mu_{0}=0$ then $x_{t}$ is a zero mean stationary process, which together with (89) means that the posterior $\pi_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}$ has a mode at the zero-path, $x_{t}=0,\quad t\geq 0$.
The parameters are set to $\Sigma_{0}=0.36$, $\phi_{0}=0.95$, $Q_{t\mid t-1}=(1-\mathinner{\!\left\lvert\phi_{0}\right\rvert}^{2})\Sigma_{0}$, $\beta=1$, $R=1$, and the mean parameter is varied according to $\mu_{0}\in\{0,0.5,1.5\}$. Thus when $\mu_{0}=0$, the state has a 10% chance of being greater than unity in magnitude. The system is for $T=2^{13}+1$ time points for each configuration.
The sub-optimal variational filter of definition˜6 is implemented and the optimization is carried out with natural gradient descent using a step-size of $1\text{\times}{10}^{-1}$ and 20 iterations per filter step. The sub-optimal filter is compared to an iterated extended Kalman filter (Bell1993), which also uses 20 iterations. The two alternatives are referred to as VF(20) and EKF(20), respectively.
Figure 4: Simulation results for the cubic sensor experiment. The rop row shows the realization of the observation sequence for each selection of $\mu_{0}$ and the bottom row shows the corresponding realization of the state sequence. The filter estimates are illustrated by a two standard deviation band around their estimates.
The results of the three simulations are shown in fig.˜4. It can be seen that EKF(20) fails to track the state completely in the case when the stationary mean of the state is zero (it gets stuck at the mode at zero). However, as the stationary mean moves away from zero VF(20) and EKF(20) become increasingly similar in performance. This is also corroborated by the sequence of root mean-square errors, which is shown in table˜1. The sum of $\chi^{2}$-statistics is shown in table˜2 from which a similar conclusion can be drawn but also that both filters are quite severely overconfident on average. In fact, the overconfidence gets worse the more time the state stays in the high signal-to-noise ratio regime.
Table 1: Root mean-square errors for VF(20) and EKF(20) in the cubic sensor problem.
It can be seen that VF(20) handles the cases when the state is seldom in the high signal-to-noise ratio regime better than EKF(20).
$\mu_{0}$ 0 0.4 0.8
VF(20) 0.411192 0.380229 0.33623
EKF(20) 0.614548 0.527413 0.410902
Table 2: Sum of $\chi^{2}$-statistics for VF(20) and EKF(20) in the cubic sensor problem. If a Gaussian filter is appropriate this sum should with high probability be $T=2^{13}+1=8193$
$\mu_{0}$ 0 0.4 0.8
VF(20) 11466.5 13863.3 20122.9
EKF(20) 8595.12 27356.0 19681.2
8 Conclusion
In this article, a complete set of variational analogue to Bayesian state estimation theory was established. In particular, the value functional associated with the backward method, $\beta_{t}$, was shown to be a lower bound on the state likelihood $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$, and the value functional associated with the forward method, $\alpha_{t}$, was shown to be a lower bound on the unnormalized filtering density $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$. Like the Bayesian state estimation recursions, the variational state estimation recursion are intractable in general. However, it has been demonstrated that they are tractable more often than their Bayesian counterparts. Additionally, a formal notion of sub-optimal variational filtering has been put forth, which ensures a time-complexity of $O(T)$. Notably this notion of variational filtering differs from that of Campbell2021. The methodology has been empirically demonstrated to be sound in a jump Gauss–Markov system with a factored approximation and in a Gauss–Markov model with log-polynomial likelihoods under a Gauss–Markov constraint. There are several issues and possibilities that can be pursued in future work. From a methodological point of view, the two most pressing issues appear to be:
  • Propositions˜1 and 2 need to compute integrals with respect to all time-marginals of the class $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$. Unfortunately, these time-marginals are analytically tractable only in a few cases such as when $\mathcal{Q}_{0\mathrel{\mathop{\ordinarycolon}}T}$ is a Gauss–Markov class.
  • The value functional recursions are intractable. ˜1 is a great start as demonstrated by its implicit use by Courts2021. However, approximation methods where the error can be reduced or even controlled could prove significant in improving state estimation performance.
A possibility of resolving the first issue is to approximate intergrals with respect to the time-marginals with Monte Carlo averages. As the value functional recursions of theorems˜3 and 4 are given in the weak sense, an immediate approach would be to resort to some Galerkin style approximation with an appropriate selection of subspace. However, these possibilities will have to explored in future work.
Acknowledgements
FT was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix A More details on jump Gauss–Markov processes
In this section, some additonal details are provided for variational state estimation in jump Gauss–Markov systems. More specifically formulae are provided for conveniently organizing the implementation. The proofs of these results reduce to rote algebra and are therefore omitted.
Lemma 4.
Let $g^{\star}(x)=\mathcal{N}(\mu^{\star},\Sigma^{*})$ and $g(x)=\mathcal{N}(\mu,\Sigma)$ then the negative relative entropy between of $g^{\star}$ with respect to $g$ is given by
[…]
Lemma 5 (Average of log-Gaussian densities).
Let $g(x\mid z)=\mathcal{N}(x;\mu(z),\Sigma(z))$ and $f(z)$ a categorical distribution. Then there is a $\bar{g}(x)=\mathcal{N}(\bar{\mu},\bar{\Sigma})$ and $\zeta$ such that
$\log\{\zeta\bar{g}(x)\}=\sum_{z}\log g(x\mid z)f(z).$ (91)
Furthermore, the parameters $\bar{\mu},\bar{\Sigma},$ and $\zeta$ are given by
$\bar{\Sigma}^{-1}$ $=\sum_{z}\Sigma^{-1}(z)f(z)$ (92a)
$\bar{\Sigma}^{-1}\bar{\mu}$ $=\sum_{z}\Sigma^{-1}(z)\mu(z)f(z)$ (92b)
$\zeta$ $=\sum_{z}\log g(\bar{\mu}\mid z)f(z)+\frac{1}{2}\log\mathinner{\lvert 2\pi\bar{\Sigma}\rvert}$. (92c)
Lemma 6 (Negative relative entropy of Gaussian conditionals).
Let $\kappa_{1}(x^{\prime}\mid x)=\mathcal{N}(x^{\prime};a_{1}(x),Q_{1})$ and $\kappa_{2}(x^{\prime}\mid x)=\mathcal{N}(x^{\prime};a_{2}(x),Q_{2})$ with $a_{i}$ an affine function with slope and intercept given by $\Phi_{i}$ and $u_{i}$, respectively for $i=1,2$. Then the negative relative entropy of $\kappa_{2}$ with respect to $\kappa_{1}$ is given by
$\int\log\Big\{\frac{\kappa_{2}(x^{\prime}\mid x)}{\kappa_{1}(x^{\prime}\mid x)}\Big\}\kappa_{1}(x^{\prime}\mid x)\operatorname{d\!}x^{\prime}=c\mathcal{N}(y;Cx,Q_{2})$ (93)
where the parameters $c$, $y$, and $C$ are given by
$\log c$ $=-\frac{1}{2}\Big[\operatorname{tr}\{Q_{2}^{-1}Q_{1}\}-\operatorname{tr}\{\operatorname{I}\}(1+\log 2\pi)-\log\mathinner{\!\left\lvert Q_{1}\right\rvert}\Big$] (94a)
$y$ $=u_{2}-u_{1}$ (94b)
$C$ $=\Phi_{1}-\Phi_{2}.$ (94c)
Appendix B More details on log-polynomial systems
In this section additional details are givne for log-polynomial systems. Some results rely on the classic formula for central moments of univariate Gaussian distributions, which is stated in lemma˜7.
Lemma 7 (Gaussian central moments).
Let $q(x)=\mathcal{N}(x;\mu,\Sigma)$ then the central moments are given by
$\int(x-\mu)^{p}q(x)\operatorname{d\!}x=\begin{cases}0,\quad p\text{ odd}\\ \Sigma^{p/2}(p-1)!!,\quad p\text{ even}\end{cases},$ (95)
where !!!! denotes the double factorial.
Lemma 8.
Let $\kappa(x\mid z)=\mathcal{N}(x;a(z),Q)$ with conditional mean function $a(x)=\Phi x+u$ and define the matrix $A_{1}(Q)$ by
$\{A_{1}(Q)\}_{n,m}$ $=\begin{cases}{n\choose m}Q^{\frac{n-m}{2}}(n-m-1)!!,\quad n-m\text{ even}\\ 0,\quad n-m\text{ odd or }m>n\end{cases}$ (96a)
$\{A_{2}(\Phi,u)\}_{n,m}$ $={n\choose m}\Phi^{m}u^{n-m}.$ (96b)
Then the following hold
$\int p(x)\kappa(x\mid z)\operatorname{d\!}x=A_{1}(Q)p(a(z))=A_{1}(Q)A_{2}(\Phi,u)p(z)$ (97)
Lemma 9.
Let $\kappa_{1}(x\mid z)=\mathcal{N}(x;a_{1}(z),Q_{1})$ and $\kappa_{2}(z\mid x)=\mathcal{N}(z;a_{2}(x),Q_{2})$ with $a_{i}$ being an affine map with slope and intercept given by $\Phi_{i}$ and $u_{i}$, respectively, for $i=1,2$. Then the negative cross-entropy of $\kappa_{2}$ relative to $\kappa_{1}$ is given by
$\begin{split}\int\log\Big\{\frac{\kappa_{2}(z\mid x)}{\kappa_{1}(x\mid z)}\Big\}\kappa_{1}(x\mid z)\operatorname{d\!}x&=-\frac{1}{2}\Big\{\log\mathinner{\!\left\lvert Q_{2}\right\rvert}-\log\mathinner{\!\left\lvert Q_{1}\right\rvert}+\operatorname{tr}(Q_{2}^{-1}\Phi_{2}Q_{1}\Phi_{2}^{*})-\operatorname{tr}(Q_{1}^{-1}Q_{1})\Big\}\\ &\quad-\frac{1}{2}\big\langle z-a_{2}\circ a_{1}(z),Q_{2}^{-1},(z-a_{2}\circ a_{1}(z))\big\rangle.\end{split}$
Additionally, in the scalar case this evaluates to a polynomial of degree two with coefficents given by
$c^{(0)}(\Phi_{1},u_{1},Q_{1})$ $=-\frac{1}{2}\Big\{\log Q_{2}-\log Q_{1}+\frac{\mathinner{\!\left\lvert\Phi_{2}\right\rvert}^{2}Q_{1}}{Q_{2}}-1+\frac{\mathinner{\!\left\lvert a_{2}(u_{1})\right\rvert}^{2}}{Q_{2}}\Big\}$ (98a)
$c^{(1)}(\Phi_{1},u_{1},Q_{1})$ $=\frac{(1-\Phi_{2}\Phi_{1})a_{2}(u_{1})}{Q_{2}}$ (98b)
$c^{(2)}(\Phi_{1},u_{1},Q_{1})$ $=-\frac{\mathinner{\!\left\lvert 1-\Phi_{2}\Phi_{1}\right\rvert}^{2}}{2Q_{2}}$ (98c)